OpenStudy (heart_offire):

Simplify. 3(2x + 4y - 2z) + 7(x + y - 4z) A) 13x + 5y - 22z B) -x - 19y + 22z C) 13x + 19y - 34z D) -x - 5y + 34z

OpenStudy (danjs):

first do the multiplication to expand out those ( ) terms, distribute the 3 and 7 into those terms

OpenStudy (danjs):

then combine the addition of the x,y,and z terms

OpenStudy (danjs):

a*(b + c + d) = a*b + a*c + a*d

OpenStudy (rose4825):

Distribute 3 and 7: (6x + 12y - 6z) + (7x + 7y - 28z)

OpenStudy (heart_offire):

so 3*2x=6x 3*4y=12y 3*2z=6z wich would make the problem (6x+12y-6z) + (7x+7y-28z)

OpenStudy (rose4825):

Now combine like terms and you will have your answer.

OpenStudy (rose4825):

13x + 19y - 34z